# AudioToolAgent: An Agentic Framework for Audio-Language Models

### v1

@@ Abstract
Large Audio-Language Models (LALMs) perform well on audio understanding tasks but lack multi-step reasoning and tool-calling found in recent Large Language Models (LLMs). This paper presents AudioToolAgent, a framework that coordinates audio-language models as tools via a central LLM agent that accesses tool adapters for audio question answering and speech-to-text. The agent selects tools, asks follow-up questions, and compares outputs for verification. Experiments with MMAU, MMAR, and MMAU-Pro show state-of-the-art accuracy: up to 74.10% on MMAU, 68.80% on MMAR, and 57.96% on MMAU-Pro. Monte Carlo sampling for shapley values across 374 configurations identifies effective agent-tool combinations. The modular design allows integration of new tools and eliminates the use of data and training costs. Code and reproduction materials are available at: https://github.com/GLJS/AudioToolAgent.
Index Terms—  Audio-Language Models, Agentic Framework, Multi-Modal Audio Understanding, Reasoning, Tool-Calling
1 Introduction
@@
While recent LALMs perform well on audio benchmarks [14, 15], few can call multiple tools [10], whereas general LLMs excel at reasoning and using external tools but lack direct audio processing. This paper combines these strengths by enabling an LLM agent to use audio models as tools. This approach achieves more adaptive audio understanding than LALMs while maintaining the deeper reasoning capabilities of general LLMs.
Fig. 1: Example of a chatbot using the AudioToolAgent framework. The agent is a large language model that coordinates the tools; the tools are audio-language models.
This paper introduces AudioToolAgent, a framework that treats audio-language models as tools and uses a central agent to coordinate them. The agent, a text-only LLM, cannot process audio directly. Instead, the LLM receives the audio file path with a question or prompt and possible answers and delegates new instructions to LALMs (tools) to be able to understand the audio. The system prompt of the LLM contains instructions for the agent on how to use the tools. Because the framework reuses pretrained state-of-the-art models, the proposed framework needs no new datasets or training. Researchers can add both new public and local tools without architectural changes.
The agent receives an audio input, a question, and answer choices. The agent uses this information to reason about the task, upon which it calls tools to be able to answer the question or prompt. For speech, the agent prioritizes speech-to-text tools to transcribe the audio. For environmental sounds or music, the agent uses general audio models to gather information. AudioToolAgent asks follow-up questions, invokes tools iteratively, compares outputs, and verifies disagreements by continuing to call tools with different inputs to increase reliability. Figure 1 shows an example of the framework in a chatbot.
Fig. 2: Schematic overview of the AudioToolAgent framework. The framework has two components: a central agent and a set of tool models. The agent is a large language model that coordinates the tools; the tools are audio-language models.
@@
    A modular architecture with an agent that coordinates audio-language tools through tool adapters. By utilizing pretrained foundational models without data or finetuning, this provides a cost-effective approach for state-of-the-art performance. Experiments on MMAU [14], MMAR [16] and MMAR-Pro [15] show that both closed-source and open-source versions of AudioToolAgent outperform prior models in several domains.
@@
    A benchmark that evaluates the effectiveness of different LLMs for reasoning on audio tasks and a benchmark that evaluates the effectiveness of different audio-language models as tools.
2 Related Work
@@
Recent advances in large language models (LLMs) have resulted in agents that perform tool calling to solve tasks [17]. This began with GPT-3.5’s function calling [18] and includes the Model Context Protocol (MCP) [19], which standardizes interactions with external tools. This work uses the ReAct framework [20]. In ReAct, the agent first reasons about the task and then performs actions to solve it. The agent selects appropriate tools to answer the question. After receiving tool responses, the agent decides whether to make additional tool calls or answer the question with the information already gathered.
@@
Recent developments include large multimodal models that integrate audio processing and agentic capabilities within a single architecture. Models such as Gemini 2.5 [5] and GPT-4o [21] handle both audio understanding and speech recognition while performing tool calling. These models can generate audio output, enabling real-time, end-to-end speech-to-speech interactions with tool use. Training these models costs substantial resources, and they remain closed-source, accessible only through API endpoints.
@@
Another relevant work, StepAudio 2 [10], received explicit training for tool calling and benchmarking on four specific tools: audio search with multimodal RAG, date and time retrieval, weather search, and web search. This model processes audio input, performs tool calls, understands audio content, and generates speech output. Training this integrated model consumed 1.356 trillion tokens over 21 days [10]. In contrast, the AudioToolAgent framework eliminates this training cost by coordinating existing pre-trained models.
@@
Similarly, other works with the same name include AudioAgent [22], which uses audio attributes to optimize prompts via a fine-tuned LLM for audio tools, and the Audio-Agent framework [23], which uses an LLM to orchestrate audio generation and editing. The current work differs by using a text-only agent that delegates audio understanding to specialized tools without fine-tuning. Instead of using fixed classifiers, AudioToolAgent queries multiple interchangeable tools and cross-checks their outputs for verification.
3 Methodology and Experimental Setup
3.1 Framework Overview
@@
The agent, a reasoning model, receives an audio file path and task description and selects tools to produce the output. It accesses audio signals only through tools. The agent identifies suitable tool calls for the task, then invokes them through structured tags: <tool_call> to initiate a request and </tool_call> to conclude it. Within these tags, the agent specifies the target tool, audio file path, and prompt. Each tool’s output enters the agent’s context, enabling it to reason and invoke additional tools as needed. The tool set includes audio understanding and speech recognition tools. Figure 2 shows a schematic visualization. To prevent runaway loops, each agent can invoke a maximum of 20 tool calls. In AudioToolAgent, the agent typically makes 5-10 calls, depending on the configuration.
@@
Instructions in the system prompt help the agent to issue follow-up tool calls when outputs conflict or remain ambiguous, and to gather targeted evidence rather than guessing. The prompt also directs the agent to resolve conflicting or ambiguous tool outputs through targeted follow-up queries and to format answers within <answer> tags for parsing. The system prompt starts as follows:
@@
System Prompt You are an expert audio analyst with access to specialized tools. Answer the question given. Put the answer between <answer> and </answer> tags. If the question is multiple choice, there is always just one choice correct. If the tool says it can’t listen to audio, try invoking the tool again. Use as many different tools as needed to answer the question, even using the same tool multiple times if needed. If initial tool outputs are conflicting or ambiguous, do not guess; instead, you must generate specific, follow-up tool calls to isolate the point of disagreement and gather more detailed evidence. The following tools are available …
@@
In this framework, all tools connect via HTTP API tool adapters for modularity. This includes public endpoints for proprietary models and self-hosted endpoints for open-source models, running on either vLLM [24] or Transformers [25]. To demonstrate the framework’s versatility, the implementation offers two configurations:
@@
  • AudioToolAgent: This configuration uses a proprietary agent and closed-source tools, accessed through public API endpoints to maximize performance. The agent is the GPT5 model [4], and the tool suite includes GPT-4o [21], Gemini 2.5 Flash [5], Voxtral [12], Qwen2.5 Omni [6], and Audio Flamingo 3 [11].
@@
  • AudioToolAgent-Open: As the primary model, this configuration uses an open-source agent with high-performing open-source audio tools to balance performance with self-hosting capabilities. The agent is DeepSeek V3.1 [1], and the tool suite includes Whisper [13], Voxtral [12], Qwen2.5 Omni [6], Audio Flamingo 3 [11], and DeSTA 2.5 [9].
3.2 Evaluation Setup
@@
The study evaluated AudioToolAgent and AudioToolAgent-Open on three benchmarks: Massive Multi-Task Audio Understanding (MMAU) [14], MMAR [16] and MMAR-Pro [15]. The MMAU benchmark includes 10,000 audio clips for multi-task audio understanding and reasoning, with 1,000 in the test-mini split and 9,000 in the test split. The experiments used only the test-mini split to reduce costs. The MMAR benchmark tests deep reasoning capabilities with 1,000 audio-question-answer triplets requiring multi-step reasoning across modalities. The MMAU-Pro benchmark measures audio intelligence using 5,304 instances (one audio example was broken) containing human expert-generated question-answer pairs across speech, sound, music, and combinations.
4 Results
@@
Table 1: Comparison of AudioToolAgent with baseline models on MMAU, MMAR, and MMAU-Pro benchmarks. AudioToolAgent achieves state-of-the-art performance across all evaluation metrics. All baseline model scores are copied from their respective original works and/or benchmark evaluations. Self-proposed, no code or API available to verify.
Table 1 summarizes the results. AudioToolAgent-Open outperforms all open-source models on average across the three benchmarks, including the individual tools it uses. AudioToolAgent outperforms most closed-source models on average, even when some models could not be used as the tools of AudioToolAgent due to no API availability.
Performance gains appear most pronounced in the Speech portions of the benchmarks (see Speech columns in Table 1). The automatic speech recognition tools in both AudioToolAgent and AudioToolAgent-Open explain this improvement. While other approaches use a single multimodal model trained for both speech recognition and audio understanding, AudioToolAgent invokes ASR models like Whisper [13] and Voxtral [12] for accurate transcription.
4.1 Ablation Study
@@
To identify the most effective configuration for AudioToolAgent, an ablation study examined 10% of the MMAU test-mini split (100 examples), analyzing agents and tools separately.
4.1.1 Agents
@@
To evaluate the LLMs capable of tool calling, the experiments used a fixed set of tools - the same tools from the open-source AudioToolAgent configuration (see Section 3.1). Each evaluation ran five independent tests per agent with different random seeds and reported the mean. This approach accounts for accuracy variations from non-deterministic inference even with fixed seeds, partly due to vendor-recommended decoding defaults such as nonzero temperature. Figure 3 shows the tested tools on the y-axis.
Figure 3 visualizes the agent ablation. Inspired by Omni-R1 [26], which showed that text-only models perform well on audio reasoning tasks, the black vertical tick on the bar plot shows each LLM’s performance without audio capabilities, which still scores well on the benchmark. The dots represent individual evaluations, with the horizontal colored bar showing the average across 5 runs.
@@
0.350.350.40.40.450.450.50.50.550.550.60.60.650.650.70.70.750.750.80.80.850.85GLM 4.5 [2]Hermes 4 405B [29]Gemini 2.5 Flash [5]Qwen 3 Next 80B A3B [3]Qwen 3 235B A22B [3]Qwen 3 Coder 480B A35B [30]GPT-5 Mini [4]GPT-5 [4]Claude Sonnet 4.0 [31]Kimi K2 [32]Deepseek V3.1 [1]Accuracy
Fig. 3: Per-model accuracies (dots; 5 runs) and mean accuracy (horizontal bar) on a subset of the MMAU test-mini split. Vertical black ticks denote the corresponding ALM without any tools.
Deepseek V3.1 outperforms all other LLMs with a mean accuracy of 0.784, followed by Kimi K2 (0.766), Claude Sonnet 4 (0.762) and GPT-5 (0.748). Based on these results, AudioToolAgent-Open uses Deepseek V3.1 and AudioToolAgent uses GPT-5. GPT-5 was chosen over Claude Sonnet 4 because it costs less in the configuration: low settings for reasoning effort and verbosity were maintained to reduce costs.
4.1.2 Tools
To quantify each tool’s contribution to system performance, this work estimates Shapley values using a two-stage, Monte Carlo approximation. In the first stage, the method generates multiple sampled permutations of the available tools. For each permutation, the system evaluates performance on the 100 examples from the MMAU subset. In the second stage, the approach calculates the final Shapley values by considering only combinations of two or more tools to determine each tool’s marginal contribution. The final Shapley value for each tool represents the average of these marginal contributions across all sampled permutations. Figure 4 displays these values. The analysis clusters the tested tools into 4 categories, shown on the y-axis of Figure 4.

### v2

@@ Abstract
Large Audio-Language Models (LALMs) perform well on audio understanding tasks but lack multistep reasoning and tool-calling found in recent Large Language Models (LLMs). This paper presents AudioToolAgent, a framework that coordinates audio-language models as tools via a central LLM agent that accesses tool adapters for audio question answering and speech-to-text. The agent reasons about which tools to invoke, how to formulate follow-up queries, and how to arbitrate conflicting tool outputs, without accessing the audio. Experiments with MMAU, MMAR, and MMAU-Pro show state-of-the-art accuracy: up to 77.50% in MMAU, 77.00% in MMAR, and 61.90% in MMAU-Pro. Monte Carlo sampling for Shapley values across 490 configurations identifies effective agent-tool combinations. The modular design allows the integration of new tools and removes the need for additional training data and training runs. The code and reproduction materials are available at https://github.com/GLJS/AudioToolAgent.
I Introduction
@@
Recent LALMs perform well on audio benchmarks [28, 13] but most lack tool use [40]. Conversely, general LLMs excel at reasoning and using external tools but lack direct audio processing. The present work combines these strengths by enabling an LLM agent to use audio models as tools.
The framework introduces AudioToolAgent, which treats audio-language models as tools and uses a central agent to coordinate them. The agent, a text-only LLM, cannot process the audio directly; it receives only the audio file path, a question or prompt, and possible answers. The agent then reasons about which tools to invoke and delegates instructions to LALMs (tools) that process the audio. Performance gains arise from two mechanisms: (i) selective use of complementary tools, combining specialized ASR with general audio understanding models, and (ii) multi-hop reasoning across multiple tools to reduce errors from individual models. Because the framework reuses pretrained models, it requires no new datasets or training.
Figure 1: Example of a chatbot using the AudioToolAgent framework. The agent is a large language model that coordinates the tools; the tools are audio-language models.
The agent receives an audio input, a question, and answer choices. The agent uses this information to reason about the task, upon which it calls tools to be able to answer the question or prompt. For speech, the agent prioritizes speech-to-text tools to transcribe the audio. For environmental sounds or music, the agent uses general audio models to gather information. AudioToolAgent asks follow-up questions, iteratively invokes tools, compares outputs, and verifies disagreements by continuing to call tools with different inputs to increase reliability. Figure 1 shows an example of the framework in a chatbot.
@@
    A reproducible, agent-based coordination framework with modular tool adapters. Using pretrained foundational models without additional data or fine-tuning, this provides a cost-effective approach to high performance on MMAU [28], MMAR [18], and MMAU-Pro [13]. Both closed-source and open-source versions of AudioToolAgent reach or exceed previously reported results on multiple benchmark subsets.
@@
    A controlled empirical study across three benchmarks with a large ablation space (agents and tools) and Shapley-based contribution analysis, establishing a measurement methodology for tool-using audio systems where tools can be swapped as new models appear without retraining.
II Related Work
@@
Recent advances in large language models (LLMs) have resulted in agents who call tools to solve tasks [24]. This started with the function calling of GPT-3.5 [23] and includes the Model Context Protocol (MCP) [3], which standardizes interactions with external tools. This work uses the ReAct framework [44]. In ReAct, the agent first reasons about the task and then performs actions to solve it. The agent selects the appropriate tools to answer the question. After receiving tool responses, the agent decides whether to make additional tool calls or answer the question with the information already gathered.
Related work in vision includes Image-of-Thought (IoT) prompting to external tools for a multimodal model to gain information [46]. The ScienceQA model [17] uses an image captioning model to provide context for a reasoning model. Similarly, SRICE [45] guides inference by allowing the model to select parts of the image through interactions with external tools.
@@
Recent developments include large multimodal models that integrate audio processing and agentic capabilities within a single architecture. Models such as Gemini 3 [10] and GPT-4o [21] handle both audio comprehension and speech recognition while performing tool calls. Training these models costs substantial resources, and they remain closed-source, accessible only through API endpoints.
@@
Another relevant work, StepAudio 2 [40], capable of audio understanding, received explicit training in tool call and benchmarking on four specific tools: audio search with multimodal RAG, date and time retrieval, weather search, and web search. This model processes audio input, performs tool calls, understands audio content, and generates speech output. Training this integrated model consumed 1.356 trillion tokens over 21 days [40]. In contrast, the AudioToolAgent framework eliminates this training cost by coordinating existing pretrained models.
@@
Similarly, other works with the same name include AudioAgent [1], which uses audio attributes to optimize prompts using a fine-tuned LLM for audio tools, and the AudioAgent framework [37], which uses an LLM to orchestrate audio generation and editing. The current work differs by using a text-only agent that delegates audio understanding to specialized tools without fine-tuning. Instead of using fixed classifiers, AudioToolAgent queries multiple interchangeable tools and cross-checks their outputs for verification.
III Methodology and Experimental Setup
III-A Framework Overview
@@
The agent, a reasoning model, receives an audio file path and task description and selects tools to produce the output. The agent accesses audio signals only through the tools. The agent identifies suitable tool calls for the task, then invokes them through structured tags: <tool_call> to initiate a request and </tool_call> to conclude it. Within these tags, the agent specifies the target tool, audio file path, and prompt. Each tool’s output enters the agent’s context, enabling it to reason and invoke additional tools as needed. The set of tools includes multiple speech recognition and audio understanding tools. Figure 2 shows a schematic visualization. To prevent runaway loops, each agent can invoke a maximum of 20 tool calls. In AudioToolAgent, the agent makes 5-10 calls, depending on the configuration.
Figure 2: Schematic overview of the AudioToolAgent framework. The framework has two components: a central agent and a set of tool models. The agent is a large language model that coordinates the tools; the tools are audio-language models.
@@
Instructions in the system prompt help the agent to issue follow-up tool calls when outputs conflict or remain ambiguous, and to gather targeted evidence. The prompt also directs the agent to resolve conflicting or ambiguous tool output through targeted follow-up queries and to format answers within <answer> tags for parsing. The system prompt starts as follows:
@@
System Prompt You are an expert audio analyst with access to specialized tools. Answer the question given. Put the answer between <answer> and </answer> tags. If the question is multiple choice, there is always just one choice correct. If the tool says it can’t listen to audio, try invoking the tool again. Use as many tools as needed to answer the question, even using the same tool multiple times if needed. If initial tool outputs are conflicting or ambiguous, do not guess; instead, you must generate specific, follow-up tool calls to isolate the point of disagreement and gather more detailed evidence. The following tools are available …
@@
In this framework, all tools connect via HTTP API tool adapters for modularity. This includes public endpoints for proprietary models and self-hosted endpoints for open-source models, running on vLLM [14] or Transformers [39]. To demonstrate the versatility of the framework, the implementation offers two configurations, each exposing four tools [^1] The exact tool registry and configuration files are available at https://github.com/GLJS/AudioToolAgent/.:
@@
  • AudioToolAgent: This configuration uses a proprietary agent and closed-source tools, accessed through public API endpoints to maximize performance. The agent is Gemini 3 Pro [10], and the tool suite includes Gemini 3 Pro [10], Qwen2.5 Omni [42], Audio Flamingo 3 [11], and Whisper [26].
@@
  • AudioToolAgent-Open: This configuration uses an open-source agent with high-performing open-source audio tools to balance performance with self-hosting capabilities. The agent is Qwen 3 235B [43], and the tool suite includes Qwen3 Omni [43], Qwen2.5 Omni [42], Audio Flamingo 3 [11], and Whisper [26].
The agent uses a multi-hop approach to reduce errors from individual tools. When the agent queries multiple tools and receives conflicting outputs, it triggers targeted re-queries: the agent formulates follow-up questions that constrain the scope to the disputed span or category (e.g., asking specifically about a contested timestamp or sound event). This iterative process continues until the evidence converges across tools or uncertainty remains explicit. By comparing answers across tools, the agent reduces reliance on any single model and mitigates biases inherent in individual tools.
III-B Evaluation Setup
@@
The study evaluated AudioToolAgent and AudioToolAgent-Open on three benchmarks: Massive Multi-Task Audio Understanding (MMAU) [28], MMAR [18] and MMAR-Pro [13]. The MMAU benchmark includes 10,000 audio clips for multi-task audio understanding and reasoning, with 1,000 in the test-mini split and 9,000 in the test split. The experiments used only the test-mini split to reduce costs. The MMAR benchmark tests deep reasoning capabilities with 1,000 audio-question-answer triplets requiring multi-step reasoning across modalities. The MMAU-Pro benchmark measures audio intelligence using 5,305 instances containing question-answer pairs generated by human experts in speech, sound, music, and combinations.
III-C Ablation Study Design
To evaluate the effectiveness of different components in the AudioToolAgent framework, an ablation study examines two aspects. First, an agent comparison benchmarks the tool-use performance of state-of-the-art reasoning LLMs while holding the tool set constant. Each agent is evaluated across five independent runs with different random seeds to account for non-deterministic inference behavior. Second, a tool contribution analysis quantifies how much each audio-language model improves performance when added to different tool combinations. The study operates on a 10% subset (100 examples) of the MMAU test-mini split to balance computational cost with statistical reliability.
For tool analysis, this work uses Shapley values [31] to quantify whether adding a tool to an existing set of tools is beneficial. Shapley values account for interactions where a tool’s value depends on which other tools are present, to be able to calculate marginal contributions in multi-tool systems. To estimate Shapley values, this work uses Monte Carlo sampling on a representative small fraction of all $2^{20}=1{,}048{,}576$ possible configurations. Without sampling, calculating all Shapley values would require evaluating every possible configuration.
The calculation consists of two stages. In the first stage, the method generates many random permutations of the available tools. For each permutation, the system calculates the accuracy on the benchmark subset of all ordered subsets (coalitions) from the empty set to the full set of tools. In the second stage, the final Shapley value for each tool is calculated across all sampled permutations. This Shapley value is the average of the tool’s marginal contribution when added to a pre-existing, non-empty set of tools. Standard errors are computed across permutations to quantify estimation uncertainty.
IV Results
@@
TABLE I: Comparison of AudioToolAgent with baseline models on MMAU, MMAR, and MMAU-Pro benchmarks. AudioToolAgent achieves state-of-the-art or near state-of-the-art performance across benchmarks. Self-proposed, no code or API available to verify.
Table I summarizes results across three benchmarks (MMAU, MMAR and MMAU-Pro). AudioToolAgent-Open achieves state-of-the-art open-source performance: 74.30% on MMAU, 65.90% on MMAR and 57.50% on MMAU-Pro, surpassing Audio Flamingo 3, Qwen2.5-Omni, and Kimi-Audio. AudioToolAgent reaches 77.50% on MMAU, 77.00% on MMAR and 61.90% on MMAU-Pro, outperforming GPT-4o Audio and Gemini 2.5 Pro.
Speech tasks show the strongest gains. AudioToolAgent achieves 78.38% on MMAU Speech and 82.99% on MMAR Speech. On MMAR speech categories, AudioToolAgent reaches 83.03% on Sound-Speech and 79.17% on Sound-Music-Speech, surpassing all baselines. This advantage comes from a specialized ASR tool (Whisper [26]), which provides value even when general omni models already support speech transcription.
ASR models trained with alignment constraints such as CTC, RNN-T, or transducer architectures return tokens only when the acoustics provides sufficient evidence, making them less prone to hallucination than general audio-language models. The ablation results in Section IV-A confirm that adding dedicated ASR tools improves accuracy on speech-heavy subsets even when omni tools are present.
Beyond speech, both models perform well on cross-domain reasoning and music. On MMAR, AudioToolAgent leads all models on Music-Speech (81.71%) and Sound-Speech (83.03%); AudioToolAgent-Open leads all open-source models on every MMAR category. Music understanding is a consistent strength: AudioToolAgent reaches 72.57% on MMAU-Pro and 72.46% on MMAU, while AudioToolAgent-Open scores 65.70% and 69.46% respectively, the highest open-source results on both benchmarks.
IV-A Ablation Study
@@
To identify the highest-accuracy configuration for AudioToolAgent, an ablation study examines 10% of the MMAU test-mini split (100 examples), analyzing agents and tools separately. The methodology for this analysis is also described in Section III-C.
IV-A1 Agents
@@
The experiments used a fixed set of five open-source tools to evaluate the LLMs capable of tool calling: Qwen2.5 Omni [42], AudioFlamingo 3 [11], DeSTA 2.5 [16], Whisper [26], and Voxtral [15]. Each evaluation ran five independent tests per agent with different random seeds and reported the mean. This approach accounts for accuracy variations from non-deterministic inference even with fixed seeds, partly due to vendor-recommended decoding defaults such as nonzero temperature. Figure 3 shows the results. Inspired by Omni-R1 [27], which showed that text-only models perform well on audio reasoning tasks, the black vertical tick shows each LLM’s no-tools baseline. The dots represent individual evaluations that show discrepancy between results; to combat this, the horizontal colored bar shows the average across 5 runs.
@@
0.350.350.40.40.450.450.50.50.550.550.60.60.650.650.70.70.750.750.80.80.850.85Mistral Large 3 [20]Deepseek V3.1 [8]GPT-5 [22]GPT-5 Mini [22]Hermes 4 405B [36]Kimi K2 [34]Gemini 2.5 Flash [7]Claude Sonnet 4.0 [2]Qwen 3 Next 80B A3B [43]Qwen 3 235B A22B [43]Gemini 3 Pro [10]Claude Opus 4.5 [2]AccuracyNo-tools LLM baselineSingle run accuracy
Figure 3: Accuracy comparison across 12 LLMs on a subset of the MMAU test-mini split. Bar is an average of 5 runs.
Claude Opus 4.5 achieves the highest mean accuracy of 0.752, followed by Gemini 3 Pro (0.734) and Qwen 3 235B (0.710). Closed-source models (Opus 4.5, Gemini 3 Pro) demonstrate the strongest orchestration capabilities, while open-source models like Qwen 3 235B and Qwen 3 Next 80B (0.700) achieve competitive results. All models benefit from tool access, with gains ranging from 15.6 percentage points (Mistral Large 3) to 31.0 percentage points (Gemini 2.5 Flash). Consistency varies: Qwen 3 235B exhibits the narrowest accuracy range (0.02), while DeepSeek V3.1 shows higher variability (range of 0.08). The gap between the weakest agent (Mistral Large 3: 0.616) and the strongest (Opus 4.5: 0.752) spans 13.6 percentage points, demonstrating that orchestration quality is a key differentiator.
Based on these results, AudioToolAgent uses Gemini 3 Pro as the agent, since Claude Opus 4.5 was too expensive. Gemini 3 Pro also serves as an audio tool. AudioToolAgent-Open uses Qwen 3 235B as the highest-performing open-source agent.
IV-A2 Tools
This work applies the Shapley values methodology (see Section III-C) to quantify each tool’s contribution to system performance. Figure 4 displays the estimated Shapley values with standard error bars computed on 490 sampled configurations across 20 tools. The analysis groups the tools into 4 categories, shown on the y-axis of Figure 4.